03 Feb 2023 2nd half ECD

Night center: 00:33 MST

Morning 12-deg twilight: 06:18 MST

Sunrise: 07:10 MST

Last night, the problems with DX adSec continued. It was used during the first half of the night (LBTI) and the start of the ECD time, but around 10:00 UT there was a fast diagnostic failure, after which DX AO operations were halted and only SX-LUCI1-AO tasks were done: optimize gain tuning, unsure about ESM testing. TCS patches were activated (PCS and OSS) and the AO software changes made. The TCS patches will remain in place; same for the AO software updates.

Tonight depends on the availability of the DX adSec. For poor weather, we have the LUCI-patrol field mapping, and if photometric, MODS standards. LBC+TMS testing and also the pointing model data collection will need to wait until Friday.

Update: Before opening, it was decided that the DX adSec could be used first half (LBTI), but not for ECD time.

Summary of the night (from Jenny; see ECD Second Half 20230203 UT Night Log for details):

Conditions were not photometric. Very thin cloud, seeing variable bouncing between 1.4-2.0”. Executed ESM testing. DX was run in parallel in seeing limited. ESM optimize gain ranges and routine was tuned. We collected comparison data with a bin1 and 2 bin4 AO reference stars. LUCI1 data with TTM was significantly improved relative to LUCI2. LUCI1 vs LUCI2 seeing limited observations were comparable but LUCI2 images were quite elongated. Tested TTM on bin2 and bin4.

Doug adjusted the Acquire Routine for TTM to include Center Star, Close Loop, Optimize Gain, Done automatically.

Wrapped up the end of the night with a light version of the AGw patrol field test after vignetting continued to plague and annoy.

Tasks to be done:

Google doc spreadsheet with outline of plan

[New, but needs DX AO]: Re-measure AGw2-LUCI2 focus offset

  • Related to IT 8807.
  • ? Have we taken any SL data with LUCI2 since the quick check made on 29 Jan. Should we take some E-time to do that?
  • Re-measuring AGw2-LUCI2 focus offset most efficiently done with AO loop closed.

ESM testing (3 hrs)

  • Testing using the ESM_V57.xml OT, testing regular ESM optimization and data quality
  • Targets M61, M67
  • Can be done in parallel with AO SX test
  • Seeing <2"

All-modes ESM (3 hrs)

  • LUCI1-ESM & LUCI2-SL
  • ESM_V57.xml M92 or M67

Check edges of LUCI-AGw patrol fields (3.5 hrs)

  • SL: both LUCI1 and LUCI2
  • poor weather
  • target prepared: NGC3190 @ RA=10, DEC=+21

[Not tonight, possibly on Friday?; bad GPS (IT8819)] Pointing Models (telescope needs to be in good thermal equilibrium)

  • collect data for a new LUCI/MODS pointing model
  • Prerequisite: GPS has to be restored, maybe Friday...
  • LUCIs
    • N375 cameras, N30_FieldStop, K-band, DIT/NDIT = 2.51s/2, FlexComp off
    • Before starting, measure rotator center using at least two track mode presets at diametrically opposed PAs
    • collimate, then collect data - provides collimation and pointing model.

Photometric standards (requires photometric conditions)

TMS (~3 hrs)

  • Prerequisite: Need to check alignment post power-outage
  • Channel-dropping on-sky with TMS updates mid-exposure:
    • selectively drop and bring back RR pairs and monitor the effect of the different configurations
    • tms_loop.py -o (-o is needed to ignore the shutter state and send corrections during exposures).
    • single filter, no dithering, series of 30-60 sec exposures, fields with fairly high star density at low & high elevation.
  • Thermal aberrations (need to analyze current data first)

Night Log

Link to google doc log: ECD Second Half 20230203 UT

-- OlgaKuhn - 02 Feb 2023
Topic revision: r7 - 03 Feb 2023, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback